you deposit 9000 for 2 years at 4% annual interest. in 2 years, you add 10000 to your accout, but the rate on your account changes to 8% annual interest (for existing balance and new deposit). leave the account untouched for additonal 20 years. how much has accumulate?
Answer #1

Answer:

Future value of $9,000 at the end of 2 years = 9,000 * (1 + 0.04)^2 = $9,734.4
New deposit + 9,734.4= 10,000 + 9,734.4 = $19,734.4
$19,734.4 earns 8% per year for 20 years.
Future value = 19,734.4 * (1 + 0.08)^20 = $91,981.19
$91,981.19 is the amount you would accumulate.